Can someone help me solve this problem?
Solve each system by the substitution method.
3x+5y=1
x=y+3

Answer by wilft1(217)

3x+5y=1
x=y+3
x is pretty much given to us, just plug it into the top system
y+3+y+3+y+3+5y=1
combine like terms
8y+9=1
subtract 9 from both sides
8y=-8
divide both sides by 8
y=-8/8
y=-1
now we have the value of y, plug that in and solve for x which will be x = 2
remember, your answer is going to be an ordered pair, (2,-1)